OpenStudy (anonymous):

what are the formulas for exponential problems? also for integrals with exponential problems

OpenStudy (anonymous):

Also for like the problems used in linear (leibnitz) equations the e power something

OpenStudy (anonymous):

by taking logs usualy @ganeshie8 would be able to tell u in detail

OpenStudy (anonymous):

alright thanks for the suggestion!

ganeshie8 (ganeshie8):

do u have any example problem ?

OpenStudy (anonymous):

yes! one second

OpenStudy (anonymous):

its from linear (leibnitz) equations. the problem was (x+1) dy/dx - y = e^3x (x+1)^2 and i got the P=-(1/1+x) and the Q=e^3x (x+1)

ganeshie8 (ganeshie8):

Yes, you want to find the integrating factor and multiply both sides ?

OpenStudy (anonymous):

yes!

ganeshie8 (ganeshie8):

\[ (x+1) y' - y = e^3x (x+1)^2 \] \[ y'+ \left(-\dfrac{1}{x+1}\right) y = e^3x (x+1) \]

OpenStudy (anonymous):

no wait! its e power 3x

ganeshie8 (ganeshie8):

\[ (x+1) y' - y = e^{3x} (x+1)^2 \] \[ y'+ \left(-\dfrac{1}{x+1}\right) y = e^{3x} (x+1) \]

ganeshie8 (ganeshie8):

IF = \(\large e^{\int -\dfrac{1}{x+1}dx} = ? \)

OpenStudy (anonymous):

ya! i got till here but then i don't know the formulas or how to solve it from here....

ganeshie8 (ganeshie8):

Alright, are you fine with below also ? IF = \(\large e^{\int -\frac{1}{x+1}dx} = e^{-\int \frac{1}{x+1}dx} = e^{- \ln |x+1|} = \dfrac{1}{x+1} \)

OpenStudy (anonymous):

i don't really know the basics, but i get the gist of it. i got till e power - integral 1/x+1 dx but then i didn't know the formulas...do you know the formulas or have a formula sheet of some sort online?

ganeshie8 (ganeshie8):

I see.... there are only two formulas you need to know

ganeshie8 (ganeshie8):

\[\large \int \frac{1}{x} dx = \ln |x| + C\] \[\large \int e^x dx = e^x + C\]

ganeshie8 (ganeshie8):

e^x is the most easy function for integrating/differentiating - as the integral or derivative of e^x is e^x itself !

ganeshie8 (ganeshie8):

you familiar with chain rule, eh ?

OpenStudy (anonymous):

not really...

ganeshie8 (ganeshie8):

Okay, there is a recommended order for doing differential equations

ganeshie8 (ganeshie8):

1) solve y' = f(x) by integrating 2) separable equations `3) linear equations`

ganeshie8 (ganeshie8):

Before doing linear equations, we should solve one/two equations of first two types ^

OpenStudy (anonymous):

do have an example of the first two types of questions you mentioned?

ganeshie8 (ganeshie8):

yeah sure :) Before doing first two types, we should know how to integrate using substitution method. let me give u one example integral : \[\large \int e^{ax+b} ~dx = ?\]

OpenStudy (anonymous):

ax+b. is that right?

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

yay! now what?

ganeshie8 (ganeshie8):

hey no, sorry ax+b is not the answer

OpenStudy (anonymous):

i have to leave now, but i'll come back online later, will you come back online here to help me again?

ganeshie8 (ganeshie8):

\[\large \int e^{ax+b} ~dx = a e^{ax+b} + C\]

ganeshie8 (ganeshie8):

Sure... I'll be around mostly :)

OpenStudy (anonymous):

how about like in 4 hours? will you be here?